OpenStudy (gabbyalicorn):

Brian has a bag that contains 14 red marbles and 12 yellow marbles. He selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Brian selects a red marble and then a yellow marble? Round your answer to the nearest percent.

OpenStudy (gabbyalicorn):

@BrainSmart

OpenStudy (gabbyalicorn):

@jabez177 @jim_thompson5910

OpenStudy (gabbyalicorn):

@Whitemonsterbunny17 @geerky42

OpenStudy (gabbyalicorn):

2/26 ?

jabez177 (jabez177):

I'm not too sure on this one...

OpenStudy (gabbyalicorn):

can you tag some other peopl please, i tagged all the people that I know that are good in math.

jabez177 (jabez177):

Except one... @Pooja195 :P

OpenStudy (gabbyalicorn):

Okay, thanks

OpenStudy (gabbyalicorn):

@Black_ninja123

OpenStudy (youngstudier):

I had this question! O_O

OpenStudy (black_ninja123):

Any answer choices?

OpenStudy (gabbyalicorn):

nope

OpenStudy (gabbyalicorn):

P(red and yellow) ≈ (blank space) %

OpenStudy (zarkon):

\[P(R_1\cap Y_2)=P(Y_2|R_1)P(R_1)\]

OpenStudy (gabbyalicorn):

?.?

OpenStudy (black_ninja123):

what is that?

OpenStudy (gabbyalicorn):

im not even that far in math to know that.. im in midddle schoool not high school..

OpenStudy (zarkon):

P(red on first and yellow on second)=P(yellow on 2nd given red on first)*P(red on first)

OpenStudy (black_ninja123):

The answer is 26% because 14/26*12/25=84/325 84/325 as a percent is 25.84%